40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 18 14:36:04 2014, den, Update, LSC, offsets in 3f and drmi stable lock on 1f prmi_rin.pdfdrmi_power.pngsrc_ol.pdf
    Reply  Sun May 18 17:57:54 2014, rana, Update, LSC, ETMX transients due to unseated bias cable ETMXbiasCableSeated.pdf
Message ID: 9969     Entry time: Sun May 18 17:57:54 2014     In reply to: 9968
Author: rana 
Type: Update 
Category: LSC 
Subject: ETMX transients due to unseated bias cable 

 The daytime crew had noticed that there were some ETMX angular shifts happening without any control or intention.

I reseated and strain relieved the bias cable coming into the backplane of the coil driver and now it seems OK.

In the 4-hour-long second trend plot below, the era before 2300 is before reseating. Afterwards, we make a couple adjustments, but so far there has been no un-asked for alignment shifts.

AdS has been run on both arms and offsets saved. Its locked on green/red and the beat frequencies are low and the amplitudes high. 

Sun May 18 23:53:37 2014: still OK...I declare it fixed.

 

Attachment 1: ETMXbiasCableSeated.pdf  254 kB  | Show | Show all
ELOG V3.1.3-